OpenStudy (anonymous):

Forgot how to do this :/ Write an equation in point-slope form for a line. a. slope 2; (4,3) b. slope -2/3; (-6,7) c. slope 0; (-4,4)

OpenStudy (anonymous):

(y-y1)=m(x-x1) that ring a bell? :) y1,x1 are your points

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Wait, would you maybe be able to do a. as an example? :/ Sorry, it'd help me though @Sheng

OpenStudy (anonymous):

np, so say #1 y-3=2(x-4) all i did was plug in slope, and coordinates simplify now y - 3 = 2x-8 y = 2x-5